OpenStudy (anonymous):

Graph the function ƒ(x) = ∛x + 2.

OpenStudy (anonymous):

Do you know how to find x intercept, y intercept?

OpenStudy (anonymous):

No, I´m new to this

OpenStudy (anonymous):

To find x-intercept, set f(x) = 0 and solve for x

OpenStudy (anonymous):

\[0 = \sqrt[3]{x} + 2\]. Can you solve it?

OpenStudy (anonymous):

No, does ^3 need to be squared?

OpenStudy (anonymous):

\[-2 = \sqrt[3]{x}\] \[(-2)^{3}= x\] x= ?

OpenStudy (anonymous):

cubed, not squared

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

Am I correct?

OpenStudy (anonymous):

Yes :) so x intercept = (-8,0) Now to find y intercept set x= 0 f(x) = 0 + 2 f(x) = 2 y intercept at (0,2)

OpenStudy (anonymous):

so the answer is x= (-8,0) and y= (0,2) ?

OpenStudy (anonymous):

No now you need to know a few points, find f(x) when x = -4,-2,-1,1,2,4 x y (aka f(x)) -3 0.5 -1 1 1 3 2 3,.2 3 3.4 Now that you have all these points, graph them and connect the dots.

OpenStudy (anonymous):

not literally connect the dot but the draw the line, yknow what I mean :)

OpenStudy (anonymous):

line/curve

OpenStudy (anonymous):

-3=3? thats how i would match them?

OpenStudy (anonymous):

remember (-8,0) and (0,2)

OpenStudy (anonymous):

no, (-3,0.5) (-1,1) (1,3) (1,3) (2, 3.2) (3, 3.4) These are points you'd put on the graph, then draw a curve/line through them.

OpenStudy (anonymous):

Ok, are there some more steps or (-3,0.5) (-1,1) (1,3) (1,3) (2, 3.2) (3, 3.4) is the answer

OpenStudy (anonymous):

I thought you have to graph it? If you graph it then draw the line then that's the answer.

OpenStudy (anonymous):

I do have to graph it but I'm new to this. Draw a line?

OpenStudy (anonymous):

get your graphing paper out if you have some, or use line paper. Draw a big cross like this